OpenStudy (anonymous):

How to find the integral of: x / ((4x+5) ^.5) dx

OpenStudy (slaaibak):

Set u = 4x + 5 x = (u-5)/4 dx = 1/4 du Now just substitute those values and the integral will be easy

OpenStudy (anonymous):

ok I got it thanks

OpenStudy (anonymous):

I though it was u substitution, but the proffesor did it with integration by parts, but I could not understant how she did it...

OpenStudy (slaaibak):

Does it look like this: \[\int\limits {x \over {\sqrt{4x+5}}}dx \] ?

OpenStudy (anonymous):

yes

OpenStudy (slaaibak):

Then he took a long/wrong route. This is an example of substitution.

OpenStudy (anonymous):

I thought so....But thank you